OpenStudy (anonymous):

Which equation illustrates the additive inverse property? Select one of the options below as your answer: A. 2 + (-2) = 0 B. 2 × 0 = 0 C. -2 + 2 = 2 + (-2) D. 1 + 2 + 3 = 1 + (2 + 3)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks could you help me on this other questions there similiar

OpenStudy (anonymous):

Which equation demonstrates the additive inverse property? Select one of the options below as your answer: A. r × 0 = r B. r + s = s + r C. r + 0 = r D. r + (-r) = 0

OpenStudy (anonymous):

It is D.